# LivePortrait: Efficient Portrait Animation with Stitching and Retargeting Control

### v1

@@ 4 Experiments
We compare our model with several non-diffusion-based methods, including FOMM [11], Face Vid2vid [5], DaGAN [6], MCNet [8], and TPSM [7], as well as diffusion-based models such as FADM [9] and AniPortrait [12]. For face vid2vid [5], we employ the implementation from [43], while for the other methods, we use the official implementations.
@@ 4.1 Self-reenactment
In Tab. 2, we present quantitative comparisons of self-reenactment. Our model slightly outperforms previous diffusion-based methods, such as FADM [9] and AniPortrait [12], in generation quality, and demonstrates better eyes motion accuracy than other methods.
@@ 4.2 Cross-reenactment
Tab. 3 shows the quantitative results of the cross-reenactment comparisons. Our model outperforms previous diffusion-based and non-diffusion-based methods in both generation quality and motion accuracy, except for the FID on TalkingHead-1KH, where the FID of the diffusion-based method AniPortrait [12] is lower than ours. The flip side is that diffusion-based methods require much more inference time than non-diffusion-based methods due to multiple denoising steps and high FLOPs. Additionally, the temporal consistency of the foreground and background is not as good compared to non-diffusion-based methods, due to the high variability of the diffusion models. This phenomenon can be observed in Fig. 6, where self-reenactment cases from testsets of VFHQ and TalkingHead-1KH are exemplified. We also illustrate the results of MegActor [17] in Fig. 6, which is also a diffusion-based method sharing a similar mutual self-attention and plugged temporal attention architecture as AniPortrait [12] and AnimateAnyone [24].
Figure 6: Temporal consistency comparisons with diffusion-based methods. These three cases are from VFHQ and TalkingHead-1KH test sets. Our animation results are in the original image space with stitching. Within the vertical circles, the statue disappears in the subsequent animated frames of FADM [9], there are pedestrian-like unnatural background movements in the animated results of AniPortrait [12], and the red banner disappears in some frames of MegActor [17]. Within the horizontal circles, there are hand-waving-like unnatural foreground movements in the animated images of both AniPortrait [12] and MegActor [17].
@@ Appendix A Benchmark Metric Details
We follow [57] to pre-process the evaluation set of the TalkingHead-1KH. In cross-reenactment, we extract 1 frame every 10 frames from each video, for a total of 24 frames as the driving sequence. For VFHQ, we extract 1 frame every 5 frames, for a total of 6 frames.

### v2

@@ 4 Experiments
We compare our model with several non-diffusion-based methods, including FOMM [11], Face Vid2vid [5], DaGAN [6], MCNet [8], and TPSM [7], as well as diffusion-based models such as FADM [9], AniPortrait [12], and X-Portrait [13]. For face vid2vid [5], we employ the implementation from [43], while for the other methods, we use the official implementations.
@@ 4.1 Self-reenactment
In Tab. 2, we present quantitative comparisons of self-reenactment. Our model slightly outperforms previous diffusion-based methods, such as FADM [9], AniPortrait [12], and X-Portrait [13], in generation quality, and demonstrates better eyes motion accuracy than other methods.
@@ 4.2 Cross-reenactment
Tab. 3 shows the quantitative results of the cross-reenactment comparisons. Our model outperforms previous diffusion-based and non-diffusion-based methods in both generation quality and motion accuracy, except for the FID on TalkingHead-1KH and the CSIM on both datasets, where the FID of the diffusion-based method AniPortrait [12] and the CSIM of X-Portrait [13] are better than ours. The flip side is that diffusion-based methods require much more inference time than non-diffusion-based methods due to multiple denoising steps and high FLOPs. Additionally, the temporal consistency of the foreground and background is not as good compared to non-diffusion-based methods, due to the high variability of the diffusion models. This phenomenon can be observed in Fig. 6, where self-reenactment cases from testsets of VFHQ and TalkingHead-1KH are exemplified. We also illustrate the results of MegActor [17] in Fig. 6, which is also a diffusion-based method sharing a similar mutual self-attention and plugged temporal attention architecture as AniPortrait [12] and AnimateAnyone [24].
Figure 6: Temporal consistency comparisons with diffusion-based methods. These three cases are from VFHQ and TalkingHead-1KH test sets. Our animation results are in the original image space with stitching. Within the vertical circles, the statue disappears in the subsequent animated frames of FADM [9], there are pedestrian-like unnatural background movements in the animated results of AniPortrait [12], and the red banner disappears in some frames of MegActor [17]. Within the horizontal circles, there are hand-waving-like unnatural foreground movements in the animated images of both AniPortrait [12] and MegActor [17], while the patterns on the clothing change in the animated images of X-Portrait [13].
@@ Appendix A Benchmark Metric Details
We follow [57] to pre-process the evaluation set of the TalkingHead-1KH. In cross-reenactment, we extract 1 frame every 10 frames from each video, for a total of 24 frames as the driving sequence. For VFHQ, we extract 1 frame every 5 frames, for a total of 6 frames. It is observed that insufficient driving video length might reduce the animation quality of X-Portrait [13]. Therefore, in the cross-reenactment experiments for X-Portrait, we use the first 250 frames of each video in the TalkingHead-1KH test set and the full videos in the VFHQ test set as the driving videos. Subsequently, we extract frames from the animated results that match the indices of the comparison frames used by other methods.